News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 9 February 2001 3.00pm EST

During the last week, Chandra completed the observing schedule as planned.

Real-time procedures were completed on 2/5/01 to re-dump SSR-B to investigate a data anomaly recorded on SSR-B, on 2/6/01 to backout the SSR-B patch and power down the recorder, and on 2/8/01 to uplink the short eclipse patch in preparation for the first eclipse on 2/11/01. In addition, routine procedures were completed to check the CPE warm-start count (remained at 1) and perform an Ephemeris update.

The planned observations for the coming week are shown below and will incorporate a real-time pass on 14 Feb to insert the LETG and on 17 Feb to perform the retraction. The schedule also includes a 35 Ks coordinated observation of UX Ari with XMM using the HRC and LETG.

------------------------------------------
1040+123                ACIS-S       Feb 12
NGC3379                 ACIS-S
2MASSJ105144+3539       ACIS-I       Feb 13
Radiation Belts
3EG J2227+6122          ACIS-I       Feb 14
Capella                 HRC-S/LETG
UX Ari                  HRC-S/LETG
Radiation Belts                      Feb 16
NGC 4593                HRC-S/LETG
HD 206267               ACIS-S/HETG  Feb 18
M31 Nucleus             ACIS-I
2MASSJ091848+2117       ACIS-I

------------------------------------------

All spacecraft subsystems continued to operate nominally.

Return to Status Reports